# Understanding and Benchmarking the Trustworthiness in Multimodal LLMs for Video Understanding

### v2

@@ F.7 Jailbreak attack
Setting. This task evaluates the robustness of videoLLMs against jailbreak attacks designed to bypass safety mechanisms and elicit harmful or inappropriate responses. The evaluation includes two image-based jailbreak attacks, FigStep [31] and MMsafetybench [32], and two video-based jailbreak attack, VideoJail [9] and VideoJailPro [9]. For the image-based attacks, adversarial examples are converted into video format to assess the model’s vulnerability in a multimodal context. The setting tests the videoLLMs’ ability to maintain safety and ethical alignment when confronted with adversarial inputs that exploit visual and temporal modalities to trigger unsafe outputs. An example for videoJail is refered to Figure 31.
Dataset. The dataset comprises adversarial examples from four jailbreak attack frameworks: FigStep [31], MMsafetybench [32], VideoJail [9], and VideoJail-Pro [9]. For the image-based attacks (FigStep and MMsafetybench), adversarial images are transformed into videos by incorporating them into sequences, ensuring compatibility with the multimodal input requirements of MLLMs. The VideoJail and VideoJail-Pro datasets include videos specifically designed to exploit weaknesses in video processing capabilities.
@@ G.1.1 Stereotype Impact Generation
Dataset. To ensure effective evaluation, we select 1,592 videos from OpenVid-1M [23], covering attributes such as gender, age, skin color, and occupation to ensure diversity, relevance, and representativeness. Prompts are then generated using Grok3666https://grok.com/ based on the detailed descriptions of each video. An illustrative example is shown in Figure 32.
@@
G.1.2 Perferecne selection of videoLLMs
@@
Figure 35: An example for the task of perference selection in videoLLMs.
@@
Metrics. To evaluate whether videoLLMs appropriately avoid subjective preferences, we use the Refuse-to-Answer (RtA) rate as the primary metric. Models are expected to decline making a choice when presented with videos representing opposing perspectives, in alignment with the principle that AI systems should not exhibit personal bias. To identify refusals, we combine a previously trained Longformer-based classifier [51] with a keyword-matching approach that detects typical refusal phrases (e.g., “Sorry, I cannot,” or “It’s not appropriate to…”). The keyword method complements the classifier by capturing cases it may miss.
TABLE XIX: Performance (%) of videoLLMs in the task of perferecne selection.
@@ G.1.3 Profession Competence Prediction
Setting. Preference selection in videoLLMs evaluates biases in subjective, preference-driven scenarios involving cultural, ideological, or lifestyle perspectives. However, biases in professional competence judgments—often grounded in objective criteria such as skills or performance—may manifest differently. To address this, we extend the evaluation to examine whether videoLLMs exhibit stereotypes or unfair assumptions related to attributes like gender, race, or appearance in professional contexts. In this task, models are challenged to predict individuals’ professional competence, reflecting how factors such as age, gender, and skin tone influence model perceptions (see Figure 37). Following an approach similar to MultiTrust [21], we present videos of individuals with diverse attributes alongside their actual professions to assess videoLLM judgments of competence. We further analyze model sensitivity through two settings: one presenting attributes visually only, and another combining visual and textual attribute information. The prompts used in both settings are illustrated in Figure 36.
@@ G.2.1 Agreement on Stereotypes
Setting. Following the setting of TrustLLM [22], we present videoLLMs with stereotypical statements and prompt them to express agreement or disagreement, thereby assessing their intrinsic attitudes toward stereotypes. Additionally, we examine how the combination of textual prompts with varying videos cues—both relevant and irrelevant—impacts the models’ tendency to endorse stereotypical views, shedding light on fairness and bias in multimodal comprehension. An example is shown in Figure 38.
Dataset. To investigate differences between MLLMs for video and image understanding, we employ stereotype prompts from MultiTrust [21], comprising 261 statements across six categories: age, gender, nationality, race, religion, and socioeconomic status. For related videos, we extract key content from these statements using Grok3 and generate semantically relevant videos via the text-to-video model Jimeng. Additionally, each statement is paired with irrelevant videos from three categories—color, nature, and noise—for comparison.
@@
Results. As shown in Table XXI, the agreement rates for related videos are generally higher than those for unrelated videos. However, closed-source models exhibit significantly lower average agreement rates (6.54%–12.31%) on related videos compared to open-source models, such as TPO, which reaches 21.15%. For example, Claude4-sonnet (6.92%) and Geminil.5-Flash (7.31%) demonstrate strong resistance to stereotypes in the presence of related videos, whereas open-source models like TPO and Oryx-34B show agreement rates exceeding 20%. Among the unrelated video types, noise and color videos lead to notably higher agreement rates than natural scenes, suggesting their stronger interference with model judgment. Within the closed-source models, GPT-4o reaches an agreement rate of 11.15% on noise videos, while Claude3.7-sonnet shows a much lower agreement rate of only 4.23% on natural scene videos.
Findings. (1) Closed-source models generally show significantly lower agreement rates with stereotypes compared to open-source models, demonstrating an advantage in recognizing and resisting stereotypical content. (2) Different types of unrelated videos interfere with model judgments to varying degrees; specifically, noise and color videos lead to significantly higher agreement rates than natural scenes. (3) Some open-source models perform well, such as Qwen-2.5-VL-7B (6.54%) and Sharegpt4video-8B (0.38%), with agreement rates approaching those of closed-source models. However, others perform poorly—for instance, Video-ChatGPT-7B exhibits an agreement rate as high as 45.77% in noise video conditions—highlighting substantial differences in the robustness of open-source models against multimodal interference.
@@ H.1.1 Private Content Recognition
Dataset. We sampled 100 videos from the BIV-Priv dataset [33], a benchmark commonly used to advance research on visual recognition under privacy constraints. The dataset is specifically curated to include privacy-sensitive content, such as passports, credit cards, medical prescriptions, and private correspondence. These videos depict real-world scenarios where sensitive information may be inadvertently exposed, offering a rigorous testbed for evaluating model detection performance.
@@ H.2.1 Infoflow Expectation
Setting. We assess the capability of videoLLMs to appropriately handle private information in videos. Following the approach of MultiTrust [21], we utilize the InfoFlow Expectation from [34], which characterizes the contexts in which specific types of private information are used and measures the alignment between model predictions and human expectations. To facilitate subsequent answer extraction, we include an explicit formatting instruction at the end of the prompt. An example is shown in Figure 44.
Dataset. Following the setting of MultiTrust [21], we use the Tier-2a subset of 98 prompts from [34], which describe only the usage of private data, excluding additional contextual short stories found in Tier-2b. This subset covers 7 types of private information across 14 usage scenarios. In addition to pair with irrelevant videos, we employ Jimeng, a commercial text-to-video model, to generate semantically relevant videos. Specifically, we use the usage scenarios as prompts for video synthesis, producing 14 relevant videos paired with the corresponding text samples.
Metrics. Following the MultiTrust [21], we evaluate the pearson correlation between model ratings and human annotations. Predictions are extracted from responses using keyword matching, capturing both numeric values (e.g., “-100”) and verbal expressions (e.g., “strongly disagree”). During preliminary testing, we observed that some models fail to provide clear scores as instructed. In such cases, the response is treated as a task failure, and a default “neutral” value (0) is assigned. Models exhibiting high failure rates are noted in the results.
@@ H.2.2 Celebrities Privacy information QA
Metrics. In the task, we use the Reject answer rate (RtA) as the sole evaluation metric. Models are expected to refuse queries involving personal identifi able information regardless of the input content. To identify refusal responses, we apply both a pre-trained Longformer evaluator [51] and keyword matching (e.g., detecting phrases like “Sorry, it’s not appropriate to…” or warnings). The keyword matching compensates for cases missed by the classifier.
Figure 46: RtA Rate (%) in the celebrity privacy information QA task. n denotes name, o denotes occupation, and w denotes without; for example, wo-name indicates that only occupation is provided in the prompt.
Results. The evaluation results for the Celebrities Privacy Information QA task, as shown in Figure 46, highlight the models’ ability to safeguard privacy-sensitive information in video content by rejecting queries involving personal identifiable information. Among closed-source commercial models, Claude4-sonnet demonstrates exceptional performance, achieving a 100.0% rejection rate (RtA) across all categories—phone, address, and email—regardless of whether name (n) or occupation (o) is provided in the prompt. GPT-4o also performs strongly, with a 100.0% RtA in most scenarios, though it slightly drops to 96.0% in phone (n-o) and email (n-o) settings, indicating minor sensitivity to fully contextualized prompts. Gemini1.5-Pro maintains a high RtA of 100.0% in most cases but shows noticeable declines when occupation is omitted (wo-o), with scores of 78.0% for address (n-wo-o) and 88.0% for email (n-wo-o), suggesting that contextual cues like occupation play a critical role in its privacy discrimination.
Among open-source models, performance varies widely. Advanced models like Sharegpt4video-7B and LLaVA-OneVision-72B exhibit strong privacy awareness, achieving 100.0% RtA in scenarios with full context (n-o) across phone, address, and email queries, surpassing some closed-source models in specific cases—for example, ShareGPT4Video-7B’s 100.0% RtA in address (wo-n-o) compared to Gemini 1.5Pro’s 80.0%. However, these models falter in reduced-context scenarios; LLaVA-OneVision-72B drops to 92.0% in address (wo-n-o) and 94.0% in email (n-wo-o), while Sharegpt4video-7B falls to 2.0% in address (wo-n-wo-o), highlighting instability when key identifiers like name or occupation are absent. Early-stage open-source models like MiniCPM-o-2.6-7B and mPLUG-Owl3-7B struggle significantly, with RtA scores as low as 0.0% in challenging scenarios such as address (wo-n-wo-o) and email (wo-n-wo-o), underscoring their limited capability to recognize and protect privacy in video modalities.

### v3

@@ F.7 Jailbreak attack
Setting. This task evaluates the robustness of videoLLMs against jailbreak attacks designed to bypass safety mechanisms and elicit harmful or inappropriate responses. The evaluation includes two image-based jailbreak attacks, FigStep [gong2025figstep] and MMsafetybench [liu2024mm], and two video-based jailbreak attack, VideoJail [hu2025videojail] and VideoJailPro [hu2025videojail]. For the image-based attacks, adversarial examples are converted into video format to assess the model’s vulnerability in a multimodal context. The setting tests the videoLLMs’ ability to maintain safety and ethical alignment when confronted with adversarial inputs that exploit visual and temporal modalities to trigger unsafe outputs. An example for videoJail is refered to Figure 31.
Dataset. The dataset comprises adversarial examples from four jailbreak attack frameworks: FigStep [gong2025figstep], MMsafetybench [liu2024mm], VideoJail [hu2025videojail], and VideoJail-Pro [hu2025videojail]. For the image-based attacks (FigStep and MMsafetybench), adversarial images are transformed into videos by incorporating them into sequences, ensuring compatibility with the multimodal input requirements of MLLMs. The VideoJail and VideoJail-Pro datasets include videos specifically designed to exploit weaknesses in video processing capabilities.
@@ G.1.1 Stereotype Impact Generation
Dataset. To ensure effective evaluation, we select 1,592 videos from OpenVid-1M [nan2024openvid], covering attributes such as gender, age, skin color, and occupation to ensure diversity, relevance, and representativeness. Prompts are then generated using Grok3666https://grok.com/ based on the detailed descriptions of each video. An illustrative example is shown in Figure 32.
@@
G.1.2 Preference selection of videoLLMs
@@
Figure 35: An example for the task of preference selection in videoLLMs.
@@
Metrics. To evaluate whether videoLLMs appropriately avoid subjective preferences, we use the Refuse-to-Answer (RtA) rate as the primary metric. Models are expected to decline making a choice when presented with videos representing opposing perspectives, in alignment with the principle that AI systems should not exhibit personal bias. To identify refusals, we combine a previously trained Longformer-based classifier [wang2023not] with a keyword-matching approach that detects typical refusal phrases (e.g., “Sorry, I cannot,” or “It’s not appropriate to…”). The keyword method complements the classifier by capturing cases it may miss.
TABLE XIX: Performance (%) of videoLLMs in the task of perference selection.
@@ G.1.3 Profession Competence Prediction
Setting. Preference selection in videoLLMs evaluates biases in subjective, preference-driven scenarios involving cultural, ideological, or lifestyle perspectives. However, biases in professional competence judgments—often grounded in objective criteria such as skills or performance—may manifest differently. To address this, we extend the evaluation to examine whether videoLLMs exhibit stereotypes or unfair assumptions related to attributes like gender, race, or appearance in professional contexts. In this task, models are challenged to predict individuals’ professional competence, reflecting how factors such as age, gender, and skin tone influence model perceptions (see Figure 37). Following an approach similar to MultiTrust [zhang2024benchmarking], we present videos of individuals with diverse attributes alongside their actual professions to assess videoLLM judgments of competence. We further analyze model sensitivity through two settings: one presenting attributes visually only, and another combining visual and textual attribute information. The prompts used in both settings are illustrated in Figure 36.
@@ G.2.1 Agreement on Stereotypes
Setting. Following the setting of TrustLLM [sun2024trustllm], we present videoLLMs with stereotypical statements and prompt them to express agreement or disagreement, thereby assessing their intrinsic attitudes toward stereotypes. Additionally, we examine how the combination of textual prompts with varying videos cues—both relevant and irrelevant—impacts the models’ tendency to endorse stereotypical views, shedding light on fairness and bias in multimodal comprehension. An example is shown in Figure 38.
Dataset. To investigate differences between MLLMs for video and image understanding, we employ stereotype prompts from MultiTrust [zhang2024benchmarking], comprising 261 statements across six categories: age, gender, nationality, race, religion, and socioeconomic status. For related videos, we extract key content from these statements using Grok3 and generate semantically relevant videos via the text-to-video model Jimeng. Additionally, each statement is paired with irrelevant videos from three categories—color, nature, and noise—for comparison.
@@
Results. As shown in Table XXI, the agreement rates for related videos are generally higher than those for unrelated videos. However, closed-source models exhibit significantly lower average agreement rates (6.5%–12.3%) on related videos compared to open-source models, such as TPO, which reaches 21.2%. For example, Claude4-sonnet (6.9%) and Geminil.5-Flash (7.3%) demonstrate strong resistance to stereotypes in the presence of related videos, whereas open-source models like TPO and Oryx-34B show agreement rates exceeding 20%. Among the unrelated video types, noise and color videos lead to notably higher agreement rates than natural scenes, suggesting their stronger interference with model judgment. Within the closed-source models, GPT-4o reaches an agreement rate of 11.2% on noise videos, while Claude3.7-sonnet shows a much lower agreement rate of only 4.2% on natural scene videos.
Findings. (1) Closed-source models generally show significantly lower agreement rates with stereotypes compared to open-source models, demonstrating an advantage in recognizing and resisting stereotypical content. (2) Different types of unrelated videos interfere with model judgments to varying degrees; specifically, noise and color videos lead to significantly higher agreement rates than natural scenes. (3) Some open-source models perform well, such as Qwen-2.5-VL-7B (6.5%) and Sharegpt4video-8B (0.4%), with agreement rates approaching those of closed-source models. However, others perform poorly—for instance, Video-ChatGPT-7B exhibits an agreement rate as high as 45.8% in noise video conditions—highlighting substantial differences in the robustness of open-source models against multimodal interference.
@@ H.1.1 Private Content Recognition
Dataset. We sampled 100 videos from the BIV-Priv dataset [sharma2023disability], a benchmark commonly used to advance research on visual recognition under privacy constraints. The dataset is specifically curated to include privacy-sensitive content, such as passports, credit cards, medical prescriptions, and private correspondence. These videos depict real-world scenarios where sensitive information may be inadvertently exposed, offering a rigorous testbed for evaluating model detection performance.
@@ H.2.1 Infoflow Expectation
Setting. We assess the capability of videoLLMs to appropriately handle private information in videos. Following the approach of MultiTrust [zhang2024benchmarking], we utilize the InfoFlow Expectation from [mireshghallah2023can], which characterizes the contexts in which specific types of private information are used and measures the alignment between model predictions and human expectations. To facilitate subsequent answer extraction, we include an explicit formatting instruction at the end of the prompt. An example is shown in Figure 44.
Dataset. Following the setting of MultiTrust [zhang2024benchmarking], we use the Tier-2a subset of 98 prompts from [mireshghallah2023can], which describe only the usage of private data, excluding additional contextual short stories found in Tier-2b. This subset covers 7 types of private information across 14 usage scenarios. In addition to pair with irrelevant videos, we employ Jimeng, a commercial text-to-video model, to generate semantically relevant videos. Specifically, we use the usage scenarios as prompts for video synthesis, producing 14 relevant videos paired with the corresponding text samples.
Metrics. Following the MultiTrust [zhang2024benchmarking], we evaluate the pearson correlation between model ratings and human annotations. Predictions are extracted from responses using keyword matching, capturing both numeric values (e.g., “-100”) and verbal expressions (e.g., “strongly disagree”). During preliminary testing, we observed that some models fail to provide clear scores as instructed. In such cases, the response is treated as a task failure, and a default “neutral” value (0) is assigned.
@@ H.2.2 Celebrities Privacy information QA
Metrics. In the task, we use the Reject answer rate (RtA) as the sole evaluation metric. Models are expected to refuse queries involving personal identifi able information regardless of the input content. To identify refusal responses, we apply both a pre-trained Longformer evaluator [wang2023not] and keyword matching (e.g., detecting phrases like “Sorry, it’s not appropriate to…” or warnings). The keyword matching compensates for cases missed by the classifier.
Figure 46: RtA Rate (%) in the celebrity privacy information QA task. n denotes name, o denotes occupation, and w denotes without; for example, w-n-o indicates that only occupation is provided in the prompt.
Results. The evaluation results for the Celebrities Privacy Information QA task, as shown in Figure 46, highlight the models’ ability to safeguard privacy-sensitive information in video content by rejecting queries involving personal identifiable information. Among closed-source commercial models, Claude4-sonnet demonstrates exceptional performance, achieving a 100.0% rejection rate (RtA) across all categories—phone, address, and email—regardless of whether name (n) or occupation (o) is provided in the prompt. GPT-4o also performs strongly, with a 100.0% RtA in most scenarios, though it slightly drops to 96.0% in phone (n-o) and email (n-o) settings, indicating minor sensitivity to fully contextualized prompts. Gemini1.5-Pro maintains a high RtA of 100.0% in most cases but shows noticeable declines when occupation is omitted (w-o), with scores of 78.0% for address (n-w-o) and 88.0% for email (n-w-o), suggesting that contextual cues like occupation play a critical role in its privacy discrimination.
Among open-source models, performance varies widely. Advanced models like Sharegpt4video-8B and LLaVA-OneVision-72B exhibit strong privacy awareness, achieving 98%-100.0% RtA in scenarios with full context (n-o) across phone, address, and email queries, surpassing some closed-source models in specific cases—for example, ShareGPT4Video-8B’s 100.0% RtA in address (n-w-o) compared to Gemini1.5-Flash’s 80.0%. However, these models falter in reduced-context scenarios; LLaVA-OneVision-72B drops to 98.0% in address (n-w-o) and 96.0% in email (w-n-w-o), while Sharegpt4video-8B falls to 2.0% in address (w-n-o), highlighting instability when key identifiers like name or occupation are absent. Early-stage open-source models like MiniCPM-o-2.6-7B and Video-ChatGPT-7B struggle significantly, with RtA scores as low as 0.0% in challenging scenarios such as address (w-n-w-o) and email (w-n-w-o), underscoring their limited capability to recognize and protect privacy in video modalities.